# Nonlinear Dynamics of Kink Configurations: From Small to Large Kink Collisions

### v1

@@ I Introduction
Topological solutions such as kinks and antikinks in (1+1)-dimensional scalar field theories are powerful tools for studying nonlinear phenomena in physics due to their stability and unique structure Rajaraman.book.1982 ; Manton.book.2004 ; Shnir.book.2018 . These configurations find applications across diverse fields, ranging from early universe cosmology Vilenkin.book.2000 to condensed matter physics and nonlinear optics Bishop.PhysD.1980 ; Optics2 . In integrable models such as the sine-Gordon model, kink-antikink scattering exhibits a simple elastic behavior accompanied by a phase shift Dmitriev.Nonlinearity.2000 ; Dmitriev.PRE.2008 ; Moradi.EPJB.2018 . However, in nonintegrable models—such as the $\phi^{4}$Rajaraman.book.1982 ; Makhankov.PhysRep.1978 ; Campbell.PhysD.1983 ; Moradi.CNSNS.2017 ; Askari.CSF.2020 ; Manton.PRL.2021 ; Almeida.EPJC.2025 , $\phi^{6}$Moradi.JHEP.2017 ; Gani.JOP.2020 ; Saadatmand.EPJB.2022 ; Adam.PRD.2022 ; Saadatmand.CSF.2024 , higher order models Khare.PRE.2014 ; Blinov.JOP.2020 ; BCM ; khare2021explicit , deformed models such as BBG ; Moradi.CSF.2022 ; Moradi.AnlPhys.2024 , and double sine-Gordon (DSG) Campbell.dsG.1986 ; Peyravi.EPJB.2009 ; Gani.EPJC.2018 ; Zolotaryuk.EPJC.2018 ; Belendryasova.JPCS.2019 ; Yerin.PRB.2021 models—a richer dynamics emerges, including the production of new kink-antikink pairs, solitary oscillations, and resonance windows. These complex behaviors have made the study of scattering a key approach to understanding nonlinear effects in physical systems.
@@
In another work introduced in Ref. Bazeia.EPJC.2017 , the investigation proposed a model derived from the deformation of the $\phi^{4}$ model, controlled by a parameter $\lambda$. This model enables a smooth transition between the sine-Gordon regime ($\lambda=0$) and vacuumless systems ($\lambda=1$), supporting two types of kink (large and small). While the work focused on the stability of solutions and braneworld applications, the scattering of kinks and antikinks in this model remained unexplored. This gap serves as the main motivation for the present study. Here, we investigate the collision of kinks and antikinks in the model for the first time, with the aim of analyzing its topological dynamics as a function of $\lambda$ and initial velocity.
Our findings reveal similarities with behaviors reported in the DSG model, notably the production of small kink-antikink pairs during large kink scattering. This overlap not only validates the proposed model, but also suggests that common mechanisms may govern such phenomena across a broader class of nonlinear systems, echoing the parameter-dependent dynamics observed in lattice models Askari.CSF.2020 ; Hadipour.PLA.2020 ; Speight.Nonlinearity.1997 ; kevrekidis.PhysicaD.2003 ; Flach.PhysRep.2008 ; Zhanna.IOPconf.2018 ; dmitriev.JPA.2005 . Moreover, the study of these effects within the framework of the model described in Ref. Bazeia.EPJC.2017 , which uniquely bridges distinct topological regimes, may offer a fresh perspective to deepen our understanding of the role of model parameters in shaping the scattering dynamics.
@@ II The Model
We explore a scalar field theory in (1+1)-dimensional spacetime, characterized by a real scalar field $\phi(x,t)$ interacting through a non-negative potential $V(\phi)$. This potential exhibits a set of minima, denoted as $\mathcal{V}=\{\phi_{1}^{(vac)},\phi_{2}^{(vac)},\phi_{3}^{(vac)},\dots\},$ that define the vacuum states of the system. The dynamics of the field are governed by the Lagrangian density
@@
The resulting equation of motion is
@@
ensuring finite energy for static configurations. Moving kinks with velocity $v$ are obtained via a Lorentz boost,
$\phi(x,t)=\phi\left(\frac{x-vt}{\sqrt{1-v^{2}}}\right).$ (8)
@@
The parameter $\lambda\in[0,1]$ tunes the model: at $\lambda=0$, we recover trigonometric functions ($\text{sn}(\chi,0)=\sin(\chi)$, $\text{cn}(\chi,0)=\cos(\chi)$, $\text{dn}(\chi,0)=1$), while at $\lambda=1$, hyperbolic functions emerge ($\text{sn}(\chi,1)=\tanh(\chi)$, $\text{cn}(\chi,1)=\text{dn}(\chi,1)=\text{sech}(\chi)$). When required, we will use $\theta=\sqrt{1-\lambda\,}$. The resulting potential is shown in Fig. 1(a),
@@
$\chi_{L,\lambda}(x)=\text{sc}^{-1}\left(\frac{\sqrt{2-\theta^{2}}}{\theta}\tanh\left(\theta\sqrt{2-\theta^{2}}x\right),\lambda\right)+2n\text{K}_{\lambda},$ (22)
where $\text{K}_{\lambda}$ is the complete elliptic integral of the first kind. For $n=0$, this solution, shown in Fig. 1(b), connects asymptotic values $\chi_{L,\lambda}(\pm\infty)=\pm\text{sc}^{-1}(\sqrt{2-\theta^{2}}/\theta,\lambda)$, with a near-origin behavior of $\chi_{L,\lambda}(x\simeq 0)\simeq(2-\theta^{2})x+\mathcal{O}(x^{2})$. As $\theta\to 0$, the asymptotic limits diverge, while for $\theta\to 1$, the solution remains well-defined. At specific $\lambda$ values, it reduces to
@@
where $R_{\lambda}=(\lambda^{2}+\lambda-2)/\sqrt{1-\lambda}$ and $r_{\lambda}=\sqrt{\lambda+1}$. These masses, plotted in Fig. 1(d), show that $M>m$ for all $\lambda\in(0,1)$. This disparity suggests that large kink collisions possess sufficient energy to produce small kink-antikink pairs across all $\lambda$. In contrast, the production of large kinks from small kink collisions may be feasible only for small $\lambda$, where $m$ and $M$ are comparable, provided the initial velocity is sufficiently high.
(a)  Deformed potential
(b)  Large kinks
(c)  Small kinks
(d)  Kink masses
Figure 1: (a) The deformed potential $U_{\lambda}(\chi)$ as defined in Eq. (20) for varying $\lambda$. (b) Spatial profiles of large kinks, $\chi_{L,\lambda}(x)$, from Eq. (22) for different $\lambda$ values. (c) Spatial profiles of small kinks, $\chi_{S,\lambda}(x)$, from Eq. (24) for different $\lambda$ values. (d) Masses of small ($m$) and large ($M$) kinks as functions of $\lambda$, calculated using Eqs. (26) and (27), respectively.
@@
for large kinks. The spectrum of Eq. (28) determines the vibrational modes of the kinks, with $\omega_{n}^{2}>0$ indicating stability. The stability potentials are illustrated in Figs. 2 and 2 for small and large kinks, respectively. For small kinks (Fig. 2), as $\lambda$ increases from 0 to 0.90, the potential wells become shallower (decreased depth) and wider (increased width), suggesting a reduction in the strength of internal modes. For large kinks (Fig. 2), the behavior is more complex: as $\lambda$ increases from 0 to 0.5, the potential wells deepen (increased depth) and narrow (decreased width), potentially enhancing vibrational modes; however, for $\lambda>0.5$, the wells become shallower and wider, weakening these modes. These variations in the stability potentials, driven by $\lambda$, directly influence the scattering dynamics, particularly the production of small kink-antikink pairs during large kink collisions.
Figure 2: (a) The stability potential $v_{S,\lambda}(x)$ for small kinks, as defined in Eq. (29), for $\lambda=0,0.25,0.50,0.75,0.90$. With increasing $\lambda$, the potential wells become shallower and wider. (b) The stability potential $v_{L,\lambda}(x)$ for large kinks, as defined in Eq. (30), for the same $\lambda$ values. For $\lambda$ from 0 to 0.54, the wells deepen and narrow, while for $\lambda>0.54$, they become shallower and wider.
@@ III.1 Small Kinks Scattering
The scattering result of small kinks is further explored in Fig. 3, which shows the final velocity $v_{f}$ as a function of the initial velocity $v_{i}$ for $\lambda=0.07$. Three distinct regimes are observed. For $v_{i}$ from 0 to 0.15, $v_{f}=0$ (red line), indicating that the kink and antikink form a bion — a bound state oscillating at the collision point with $n$ bounces. For $v_{i}$ between 0.15 and 0.53 (purple line), the small kinks separate after collision, with $v_{f}$ increasing almost linearly with $v_{i}$, although $v_{f}<v_{i}$ due to energy losses in vibrational modes and radiation. For $v_{i}>0.53$ (blue line), the collision results in the creation of a pair of large kinks. Since the mass of large kinks ($M$) is greater than that of small kinks ($m$), the final velocity of large kinks is lower than that of small kinks for the same $v_{i}$, consistent with energy conservation.
@@
Figure 3: Critical velocity, vibrational mode, and scattering outcomes for small kinks. (a) The critical velocity $v_{c}$ as a function of $\lambda$, showing a peak at $\lambda=0.54$. (b) The first vibrational mode $\omega_{1}^{2}$ of small kinks as a function of $\lambda$, decreasing from 3.938 at $\lambda=0.09$ to $0.067$ at $\lambda=0.99$. (c) Final velocity $v_{f}$ versus initial velocity $v_{i}$ for $\lambda=0.07$, showing bion formation ($v_{i}<0.15$), small kink separation ($0.15<v_{i}<0.53$), and large kink creation ($v_{i}>0.53$).
@@
To illustrate the scattering dynamics in more detail, Fig. 4 shows the evolution of the field $\chi(x,t)$ for $\lambda=0.07$ at different initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$, for kink $(k)$ and antikink $(ak)$. In Fig. 4(a), for $v_{0k}=-v_{0ak}=0.1$, the kink and antikink become trapped after collision, forming a bion that oscillates at $x\approx 0$ with $n$ bounces. This is consistent with Fig. 3, as $v_{i}=0.1$ is below the critical velocity $v_{c}=0.15$ for $\lambda=0.07$. In Fig. 4(b), with $v_{0k}=-v_{0ak}=0.2$, the kink and antikink separate after collision, moving away symmetrically, as $v_{i}=0.2$ exceeds $v_{c}$. Similarly, in Fig. 4(c), for $v_{0k}=-v_{0ak}=0.5$, the small kinks separate with a higher final velocity, still within the small kink regime ($v_{i}<0.53$). Finally, in Fig. 4(d), for $v_{0k}=-v_{0ak}=0.6$, the collision leads to the creation of large kinks, consistent with the blue line in Fig. 3. The large kinks, having a greater mass, move with a lower final velocity compared to small kinks at the same initial velocity.
(a) $v_{0k}=-v_{0ak}=0.1$             kink and antikink trapped
(b) $v_{0k}=-v_{0ak}=0.2$
(c) $v_{0k}=-v_{0ak}=0.5$
(d) $v_{0k}=-v_{0ak}=0.6$          creation of large kinks
@@
Figure 4: Small kinks scattering for $\lambda=0.07$ and different values of initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$ for all cases. (a) At $v_{i}=0.1$, the kink and antikink form a bion. (b) At $v_{i}=0.2$, the kink and antikink separate after collision. (c) At $v_{i}=0.5$, separation occurs with a higher final velocity. (d) At $v_{i}=0.6$, large kinks are created.
@@
The resonance phenomena in small kink-antikink collisions are explored in Fig. 5, which examines the number of bounces and final velocities for $\lambda=0.25$, 0.5, and 0.75. In Fig. 5(a), for $\lambda=0.25$, the number of bounces is shown as a function of $v_{i}$. Below $v_{i}=0.324784$, the kinks form a bion, exhibiting $n$ bounces as an oscillatory bound state. For $v_{i}$ from 0.324784 to the critical velocity $v_{c}=0.342718$, resonance windows are observed, with the number of bounces reaching $2$ or more, indicating that the kinks collide multiple times before separating due to energy exchange with their vibrational modes. The width of this resonance window is $0.017934$. For $v_{i}>0.342718$, the number of bounces is typically $1$, as the kinks separate after a single collision. The corresponding final velocities are shown in Fig. 5(b). For $v_{i}$ from $0.324784$ to $0.342718$, where two-bounce or higher cases occur, $v_{f}$ varies between 0 and $0.20$, reflecting the complex energy exchange in vibrational modes and radiation. Notably, in cases where $v_{f}$ reaches higher values (e.g., up to $0.20$), energy is transferred from the internal vibrational modes of the kinks to their translational modes, leading to a higher final velocity. For $v_{i}>0.342718$, where one-bounce cases dominate, $v_{f}$ increases with $v_{i}$, but remains lower than the maximum values observed in the two-bounce regime, as less energy is exchanged in multiple collisions. This indicates that the final velocity in the two-bounce regime can be higher than in the one-bounce regime, due to the resonant energy transfer from internal to translational modes in the former case.
@@
Figure 5: Small kink-antikink collision, (a) and (b) correspond to simulations with parameter $\lambda=0.25$, (c) and (d) correspond to $\lambda=0.5$, while (e) and (f) correspond to $\lambda=0.75$. (a), (c), and (e) depict the number of bounces as a function of initial velocity. Finally, (b), (d), and (f) present the final velocity for cases with one and two bounces as a function of initial velocity, $X_{0k}=-10$.
The observed trends in resonance phenomena are directly related to the vibrational mode of small kinks. As $\lambda$ increases, the frequency $\omega_{1}^{2}$ decreases (Fig. 3), reducing the ability of small kinks to store energy in their vibrational modes. Consequently, resonance windows become less prominent at higher $\lambda$ values, leading to simpler scattering dynamics. The width of the resonance windows, however, shows a non-monotonic trend: it increases from 0.017934 at $\lambda=0.25$ to 0.045884 at $\lambda=0.5$, but then slightly increases further to 0.051614 at $\lambda=0.75$, despite the near absence of resonance windows at this $\lambda$. The critical velocity $v_{c}$ increases from 0.342718 at $\lambda=0.25$ to 0.374851 at $\lambda=0.5$, reflecting the increasing difficulty of separation due to changes in the potential $U_{\lambda}(\chi)$. However, $v_{c}$ slightly decreases to 0.370429 at $\lambda=0.75$, consistent with the trend in Fig. 3, possibly due to the weakening of vibrational modes. Additionally, the absence of vibrational modes in large kinks, as determined earlier, explains why large kink creation is only observed at low $\lambda$ values (e.g., $\lambda=0.07$), where the mass difference MmM-m is smaller (Fig. 1(d)), making the transition energetically feasible.

### v2

@@ I Introduction
Topological solutions such as kinks and antikinks in (1+1)-dimensional scalar field theories are powerful tools for studying nonlinear phenomena in physics due to their stability and unique structure Rajaraman.book.1982 ; Vilenkin.book.2000 ; Manton.book.2004 ; Shnir.book.2018 . These configurations find applications across diverse fields, ranging from early universe cosmology Vilenkin.book.2000 to condensed matter physics and nonlinear optics Bishop.PhysD.1980 ; Optics2 . In integrable models such as the sine-Gordon model, kink-antikink scattering exhibits simple elastic behavior accompanied by a phase shift Dmitriev.Nonlinearity.2000 ; Dmitriev.PRE.2008 ; Moradi.EPJB.2018 . However, in nonintegrable models—such as the $\phi^{4}$Rajaraman.book.1982 ; Makhankov.PhysRep.1978 ; Campbell.PhysD.1983 ; Moradi.CNSNS.2017 ; Askari.CSF.2020 ; Manton.PRL.2021 ; Almeida.EPJC.2025 , $\phi^{6}$Moradi.JHEP.2017 ; Gani.JOP.2020 ; Saadatmand.EPJB.2022 ; Adam.PRD.2022 ; Saadatmand.CSF.2024 , higher order models Khare.PRE.2014 ; Blinov.JOP.2020 ; BCM ; khare2021explicit , deformed models such as BBG ; Moradi.CSF.2022 ; Moradi.AnlPhys.2024 , and double sine-Gordon (DSG) Campbell.dsG.1986 ; Peyravi.EPJB.2009 ; Gani.EPJC.2018 ; Zolotaryuk.EPJC.2018 ; Belendryasova.JPCS.2019 ; Yerin.PRB.2021 models—a richer dynamics emerges, including the production of new kink-antikink pairs, solitary oscillations, and resonance windows. These complex behaviors have made the study of scattering a key approach to understanding nonlinear effects in physical systems.
@@
In another work introduced in Ref. Bazeia.EPJC.2017 , the investigation proposed a model derived from the deformation of the $\phi^{4}$ model, controlled by a parameter $\lambda$. This model enables a smooth transition between the sine-Gordon regime ($\lambda=0$) and vacuumless systems ($\lambda=1$), supporting two types of kinks (large and small). While the work focused on the stability of solutions and braneworld applications, the scattering of kinks and antikinks in this model remained unexplored. This gap serves as the main motivation for the present study. Here, we investigate the collision of kinks and antikinks in the model for the first time, with the aim of analyzing its topological dynamics as a function of $\lambda$ and initial velocity.
Our findings reveal similarities with behaviors reported in the DSG model, notably the production of small kink-antikink pairs during large kink scattering. This overlap not only validates the proposed model but also suggests that common mechanisms may govern such phenomena across a broader class of nonlinear systems, echoing the parameter-dependent dynamics observed in lattice models Askari.CSF.2020 ; Hadipour.PLA.2020 ; Speight.Nonlinearity.1997 ; kevrekidis.PhysicaD.2003 ; Flach.PhysRep.2008 ; Zhanna.IOPconf.2018 ; dmitriev.JPA.2005 . Moreover, the study of these effects within the framework of the model described in Ref. Bazeia.EPJC.2017 , which uniquely bridges distinct topological regimes, may offer a fresh perspective to deepen our understanding of the role of model parameters in shaping the scattering dynamics.
@@ II The Model
We explore a scalar field theory in (1+1)-dimensional spacetime, characterized by a real scalar field $\phi(x,t)$ interacting through a non-negative potential $V(\phi)$. This potential exhibits a set of minima, denoted as $\mathcal{V}=\{\phi_{1}^{(vac)},\phi_{2}^{(vac)},\phi_{3}^{(vac)},\dots\},$ that define the vacuum states of the system. The dynamics of the field are governed by the Lagrangian density Rajaraman.book.1982 ; Vilenkin.book.2000 ; Manton.book.2004 ; Shnir.book.2018
@@
The resulting equation of motion is Rajaraman.book.1982 ; Vilenkin.book.2000 ; Manton.book.2004 ; Shnir.book.2018
@@
ensuring finite energy for static configurations. Moving kinks with velocity $v$ are obtained via a Lorentz boost $x\to\gamma(x-vt)$ with the Lorentz factor $\gamma=1/\sqrt{1-v^{2}}$ Rajaraman.book.1982 ; Vilenkin.book.2000 ; Manton.book.2004 ; Shnir.book.2018 . In this case, the field configuration changes as follows
$\phi(x,t)\to\phi\left(\frac{x-vt}{\sqrt{1-v^{2}}}\right).$ (8)
@@
The parameter $\lambda\in[0,1]$ tunes the model: at $\lambda=0$, we recover trigonometric functions ($\text{sn}(\chi,0)=\sin(\chi)$, $\text{cn}(\chi,0)=\cos(\chi)$, $\text{dn}(\chi,0)=1$), while at $\lambda=1$, hyperbolic functions emerge ($\text{sn}(\chi,1)=\tanh(\chi)$, $\text{cn}(\chi,1)=\text{dn}(\chi,1)=\text{sech}(\chi)$). When required, we will use $\theta=\sqrt{1-\lambda\,}$. The resulting potential is shown in Fig. 1,
@@
$\chi_{L,\lambda}(x)\!=\text{sc}^{-1}\left(\frac{\sqrt{2-\theta^{2}}}{\theta}\tanh\left(\theta\sqrt{2-\theta^{2}}x\right),\lambda\right)\!+2n\text{K}_{\lambda}.$ (22)
Here $\text{K}_{\lambda}$ is the complete elliptic integral of the first kind. See Ref. Bazeia.EPJC.2017 for other details on this issue. For $n=0$, the above solution, shown in Fig. 1, connects asymptotic values $\chi_{L,\lambda}(\pm\infty)=\pm\text{sc}^{-1}(\sqrt{2-\theta^{2}}/\theta,\lambda)$, with a near-origin behavior of $\chi_{L,\lambda}(x\simeq 0)\simeq(2-\theta^{2})x+\mathcal{O}(x^{2})$. As $\theta\to 0$, the asymptotic limits diverge, while for $\theta\to 1$, the solution remains well-defined. At specific $\lambda$ values, it reduces to
@@
where $R_{\lambda}=(\lambda^{2}+\lambda-2)/\sqrt{1-\lambda}$ and $r_{\lambda}=\sqrt{\lambda+1}$. These masses, plotted in Fig. 1, show that $M>m$ for all $\lambda\in(0,1)$. This disparity suggests that large kink collisions possess sufficient energy to produce small kink-antikink pairs across all $\lambda$. In contrast, the production of large kinks from small kink collisions may be feasible only for small $\lambda$, where $m$ and $M$ are comparable, provided the initial velocity is sufficiently high.
Figure 1: (a) The deformed potential $U_{\lambda}(\chi)$ as defined in Eq. (20) for different values of $\lambda$. (b) Spatial profiles of large kinks, $\chi_{L,\lambda}(x)$, from Eq. (22) for different values of $\lambda$. (c) Spatial profiles of small kinks, $\chi_{S,\lambda}(x)$, from Eq. (24) for different values of $\lambda$. (d) Masses of small ($m$) and large ($M$) kinks as functions of $\lambda$, calculated using Eqs. (26) and (27), respectively.
@@
for large kinks. The spectrum of Eq. (28) determines the vibrational modes of the kinks, with $\omega_{n}^{2}>0$ indicating stability. The stability potentials are illustrated in Figs. 2 and 2 for small and large kinks, respectively. For small kinks (Fig. 2), as $\lambda$ increases from 0 to 0.90, the potential wells become shallower (decreased depth) and wider (increased width), suggesting a reduction in the strength of internal modes. For large kinks (Fig. 2), the behavior is more complex: as $\lambda$ increases from 0 to 0.5, the potential wells deepen (increased depth) and narrow (decreased width), potentially enhancing vibrational modes; however, for $\lambda>0.5$, the wells become shallower and wider, weakening these modes. These variations in the stability potentials, driven by $\lambda$, directly influence the scattering dynamics, particularly the production of small kink-antikink pairs during large kink collisions. We further note that the stability potential for large kinks depicted in Fig. 2 shows an interesting behavior, a change in its profile, from modified Poeschl-Teller to the volcano shape as $\lambda$ increases in the interval $\lambda\in(0,1)$. See Ref. Bazeia.EPJC.2017 for other details on this issue.
Figure 2: (a) The stability potential $v_{S,\lambda}(x)$ for small kinks, as defined in Eq. (29), for $\lambda=0,0.25,0.50,0.75,0.90$. One sees that increasing $\lambda$, the potential well which has the Poeschl-Teller shape, becomes shallower and wider. (b) The stability potential $v_{L,\lambda}(x)$ for large kinks, as defined in Eq. (30), for the same $\lambda$ values. For $\lambda$ from 0 to 0.54, the wells deepen and narrow, while for $\lambda>0.54$, they become shallower and wider. We also note a change in the shape of the potential, from the reflectionless modified Poeschl-Teller to the volcano type as $\lambda$ increases towards unit.
@@ III.1 Small Kinks Scattering
The scattering of small kinks is further explored in Fig. 3, which shows the final velocity $v_{f}$ as a function of the initial velocity $v_{i}$ for $\lambda=0.07$. Three distinct regimes are observed. For $v_{i}$ from 0 to 0.15, $v_{f}=0$ (red line), the result indicates that the kink and antikink form a bion — a bound state oscillating at the collision point with $n$ bounces. For $v_{i}$ between 0.15 and 0.53 (purple line), the small kinks separate after the collision, with $v_{f}$ increasing almost linearly with $v_{i}$, although $v_{f}<v_{i}$ occurs due to the energy loss in vibrational modes and radiation. For $v_{i}>0.53$ (blue line), the collision results in the creation of a pair of large kinks. Since the mass of large kinks ($M$) is greater than that of small kinks ($m$), the final velocity of large kinks is lower than that of small kinks for the same $v_{i}$, consistent with energy conservation.
@@
Figure 3: (a) The critical velocity $v_{c}$ as a function of $\lambda$, showing a peak at $\lambda=0.54$. (b) The first vibrational mode $\omega_{1}^{2}$ of small kinks as a function of $\lambda$, decreasing from 3.938 at $\lambda=0.09$ to $0.067$ at $\lambda=0.99$. (c) Final velocity $v_{f}$ versus initial velocity $v_{i}$ for $\lambda=0.07$, showing bion formation ($v_{i}<0.15$), small kink separation ($0.15<v_{i}<0.53$), and large kink creation ($v_{i}>0.53$).
@@
To illustrate the scattering dynamics in more detail, Fig. 4 shows the evolution of the field $\chi(x,t)$ for $\lambda=0.07$ at different initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$, for kink $(k)$ and antikink $(ak)$. In Fig. 4, for $v_{0k}=-v_{0ak}=0.1$, the kink and antikink become trapped after collision, forming a bion that oscillates at $x\approx 0$ with $n$ bounces. This is consistent with Fig. 3, as $v_{i}=0.1$ is below the critical velocity $v_{c}=0.15$ for $\lambda=0.07$. In Fig. 4, with $v_{0k}=-v_{0ak}=0.2$, the kink and antikink separate after collision, moving away symmetrically, as $v_{i}=0.2$ exceeds $v_{c}$. Similarly, in Fig. 4, for $v_{0k}=-v_{0ak}=0.5$, the small kinks separate with a higher final velocity, still within the small kink regime ($v_{i}<0.53$). Finally, in Fig. 4, for $v_{0k}=-v_{0ak}=0.6$, the collision leads to the creation of large kinks, consistent with the blue line in Fig. 3. The large kinks, having greater mass, move with a lower final velocity compared to small kinks at the same initial velocity.
@@
Figure 4: Illustration of the small kinks scattering for $\lambda=0.07$ and different values of initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$ for all cases. (a) At $v_{i}=0.1$, the kink and antikink form a bion. (b) At $v_{i}=0.2$, the kink and antikink separate after collision. (c) At $v_{i}=0.5$, separation occurs with a higher final velocity. (d) At $v_{i}=0.6$, large kinks are created.
@@
The resonance phenomena in small kink-antikink collisions are explored in Fig. 5, which examines the number of bounces and final velocities for $\lambda=0.25$, 0.5, and 0.75. In Fig. 5(a), for $\lambda=0.25$, the number of bounces is shown as a function of $v_{i}$. Below $v_{i}=0.324784$, the kinks form a bion, exhibiting $n$ bounces as an oscillatory bound state. For $v_{i}$ from 0.324784 to the critical velocity $v_{c}=0.342718$, resonance windows are observed, with the number of bounces reaching $2$ or more, indicating that the kinks collide multiple times before separating due to energy exchange with their vibrational modes. The width of this resonance window is $0.017934$. For $v_{i}>0.342718$, the number of bounces is typically $1$, as the kinks separate after a single collision. The corresponding final velocities are shown in Fig. 5(b). For $v_{i}$ from $0.324784$ to $0.342718$, where two-bounce or higher cases occur, $v_{f}$ varies between 0 and $0.20$, reflecting the complex energy exchange in vibrational modes and radiation. Notably, in cases where $v_{f}$ reaches higher values (e.g., up to $0.20$), energy is transferred from the internal vibrational modes of the kinks to their translational modes, leading to a higher final velocity. For $v_{i}>0.342718$, where one-bounce cases dominate, $v_{f}$ increases with $v_{i}$, but remains lower than the maximum values observed in the two-bounce regime, as less energy is exchanged in multiple collisions. This is an interesting result, since it indicates that the final velocity in the two-bounce regime can be higher than in the one-bounce regime, due to the resonant energy transfer from internal to translational modes in the former case.
@@
Figure 5: Small kink-antikink collision, (a) and (b) correspond to simulations with parameter $\lambda=0.25$, (c) and (d) correspond to $\lambda=0.5$, while (e) and (f) correspond to $\lambda=0.75$. (a), (c), and (e) depict the number of bounces as a function of initial velocity and (b), (d), and (f) present the final velocity for cases with one and two bounces as a function of initial velocity. We are using $X_{0k}=-10$.
The observed trends in resonance phenomena are directly related to the vibrational mode of small kinks. As $\lambda$ increases, the frequency $\omega_{1}^{2}$ decreases (Fig. 3), reducing the ability of small kinks to store energy in their vibrational modes. Consequently, resonance windows become less prominent at higher $\lambda$ values, leading to simpler scattering dynamics. The width of the resonance windows, however, shows a non-monotonic trend: it increases from 0.017934 at $\lambda=0.25$ to 0.045884 at $\lambda=0.5$, but then slightly increases further to 0.051614 at $\lambda=0.75$, despite the near absence of resonance windows at this $\lambda$. The critical velocity $v_{c}$ increases from 0.342718 at $\lambda=0.25$ to 0.374851 at $\lambda=0.5$, reflecting the increasing difficulty of separation due to changes in the potential $U_{\lambda}(\chi)$. However, $v_{c}$ slightly decreases to 0.370429 at $\lambda=0.75$, consistent with the trend in Fig. 3, possibly due to the weakening of vibrational modes. Additionally, the absence of vibrational modes in large kinks, as determined earlier, explains why large kink creation is only observed at low $\lambda$ values (e.g., $\lambda=0.07$), where the mass difference MmM-m is smaller (Fig. 1), making the transition energetically feasible.